Status Report

NASA Space Station On-Orbit Status 5 Dec 2003

By SpaceRef Editor
December 6, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 5 Dec 2003

All ISS systems continue to function nominally, except as noted previously or below.

FE Alexander Kaleri conducted the second session of the current Russian TEKh-20 Plasma Crystal-3 (PK-3) experiment series.  [After wake-up, Sasha again activated the turbopump in the Service Module’s Transfer Compartment (SM PkhO), reported to MCC-M and started the evacuation of the vacuum chamber (ZB) in the SM Work Compartment (RO), then monitored Part 2 of PK-3 operations.  At Experiment Start + 23 min., live TV downlink via S-band provided coverage for ~7 min.  The turbopump was to be deactivated tonight at ~4:25pm EST.  The experiment is performed on plasma, i.e., fine particles charged and excited by HF (high frequency) radio power inside the evacuated work chamber.  Main objective today was to obtain (on video cassette) homogenous dust textures, i.e., without voids in the center, at set (preprogrammed) HF discharge powers, pressures, and various numbers of particles.  The experiment is conducted in automated mode.]

Kaleri worked on the SM water supply system, pressurizing the collapsed bladders of the Rodnik water storage tanks (BV1 & BV2) in the 12P/Progress-248 vehicle, a 1.5 hr. task, as leak-checkup preparatory to the regular transfer of liquid waste (urine) from EDV-U containers to the Progress for disposal.  [Each of the two spherical Rodnik tanks consists of a hard shell with a soft membrane (bladder) composed of elastic fluoroplastic, which is used to expel water from the tank by compressed air pumped into the tank volume surrounding the bladder.]

CDR Michael Foale completed preparations for resuming the PFMI experiment (Pore Formation and Mobility Investigation) next week, installing and tuning the PFMI hardware in the MSG (Microgravity Science Glovebox) and setting up the MSG video equipment.

At 10:20am, the Science Officer held a 25-min. teleconference with the Principal Investigator and experiment developers of CBOSS-FDI (Cellular Biotechnology Operations Support System/Fluid Dynamics Investigation), the resumption of which is scheduled in the near future.   [The discussion concerned technical and operational issues of the upcoming FDI sessions, in particular the evaluation results of the recent photography setup ops which experimented with optimum distinguishability of TCM (tissue culture module) contents with (a) the flash shielded by a paper tube vs. (b) shooting the transparent TCMs in a middeck locker environment.]

Foale also worked on the structural loads data collected with strain gauges and accelerometers during yesterday’s TVIS treadmill testing, downloading them from IWIS NCU (internal wireless instrumentation system/network control unit) #1033 and the three RSU/remote sensing units #1026, #1027, and #1028 to a hard-disk card for subsequent OCA downlink to the ground.

In preparation for the internal inspection of the TVIS gyroscope scheduled for Tuesday and Wednesday next week, Mike tagged up with ground specialists at 10:45am to discuss procedures for the IFM (in-flight maintenance).  [A 29-page preliminary procedures draft was uplinked last night, covering the extensive job of removing TVIS from the Service Module (SM) “pit”, resetting the roll-gyro-enabling DIP switch so pitch and yaw stabilizers can be turned on, removing and opening the gyro for a closer look, and photographing its internal components to help the ground determine cause & extent of the failure so that proper replacements can be manifested on Progress 13P (pending Russian concurrence).]

FE Kaleri worked briefly in the Soyuz TMA-3, placing an identification label on a ZUP-86 safety device of the Soyuz’ thermal control system.

Completing the microbial analysis of water samples collected on 12/3 from all three SM potable water ports with the WMK (water monitoring kit), SO Foale collected the microbiological data of the incubated MCDs (microbial capture devices), including visual analysis for colony growth, and recorded the data on the MEC (medical equipment computer) for later ground evaluation.

Sasha performed the regular once-a-week maintenance reboot on the operational PCS laptops.  He also restarted the OCA comm router laptop (every two weeks).

The crew worked out on TVIS, CEVIS cycle, RED exerciser and VELO bike.

The CDR did another data files transfer from the TVIS/CEVIS/RED exercise equipment to the MEC via memory card and RED log entries, for downlink on OCA comm.

Later, Mike performed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleted them on the HRM.  

Foale also terminated the 24-hour discharging process for spacesuit (EMU) batteries #2032 & #2033 in the Airlock (A/L)’s battery stowage assembly (BSA), then stowed them in the A/L after resetting their 50-day clock.

Kaleri used the CDMK (carbon dioxide monitoring kit) to take a reading of the cabin air’s ad hoc CO2 partial pressure, for calldown to the ground (along with the battery status) for use in trending analyses.  The CDMK was then restowed at the SM Central Post.

Sasha also completed the daily SOZh life support systems maintenance (including toilet facility, food containers, water containers and solid waste containers) and prepared the daily IMS (inventory management system) update file, while Mike performed the regular routine status checkup of the autonomous Increment 8 payloads.

The FE held his regular 15-min. tag-up with ground specialists via S-band to discuss the weekly IMS (inventory management system) status update and open storage/location issues.

At 9:50am EST, the crew engaged in a ham radio pass with amateur radio fans at the Berufliches Schulzentrum für Elektrotechnik (Vocational School Center for Electrical Engineering) in Dresden, Germany.  [Questions by students were uplinked beforehand.  The Berufliches Schulzentrum für Elektrotechnik, with more than 2500 students and 100 teachers, is a school for various educational and vocational fields.  It has an amateur radio club, founded in 1998, with the main purpose of preparing pupils in license courses for an amateur radio exam.  Many students obtained their amateur license in preparation for today’s contact with the ISS crew.]

At 2:00pm, the crew was scheduled to have their weekly teleconference with the ISS Flight Director at MCC-H, followed by the weekly conference with the JSC Astronaut Office (Kent Rominger).

At 10:10am EST tomorrow, ISS attitude will be changed from the current LVLH (local vertical/local horizontal) earth orientation to XPOP (x-axis perpendicular to orbit plane), facing the sun always with the same side as the station revolves around Earth.   [The maneuver will be conducted by the Russian MCS (motion control system) and thrusters, with the U.S. CMGs standing by for subsequent attitude hold if needed.  All maneuver steps were designed to minimize the CMG gimbal rates (i.e., stress).  PPLs (preplanned loads) were put in place to ensure that 5Hz vibratory data of all three gyros are captured and that data stored on the ZOE (zone-of-exclusion) recorder are downlinked for all LOS (loss-of-signal) periods occurring during the three hours following initialization of the CMG’s momentum manager.]

Tonight at 6:00pm, facility/ground systems specialists at MCC-H will “swing” (go) to LSOS (limited station operations support) mode for a period of two hours, for independent verification testing (IVT) of this mode, involving JSC’s Building 30. [During this time, the ground will have no ISS voice, telemetry, or command uplink.  The crew will be “prime” for responding to Caution alerts from the P6 solar arrays (BGA 2B & 4B motor trip and stall condition alerts).  They were reminded to turn up the respective signal volume before going to sleep.  A checkout of the LSOS mode to verify proper config before the actual swing from nominal operations to LSOS was performed last night at 6:00pm.]

Next week (12/19), TsUP/Moscow will observe the 30th
Anniversary of its Main Real-Time Mission Control Group, with a celebration at the M.I. Kalinin Cultural Center in Korolev in the Moscow suburbs.  A congratulatory message from the ISS crew to be replayed at the event from tape is scheduled for tomorrow.

Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, were Haze event, India (Dynamic event. The haze event continues, with visibilities reduced to 1-2 miles in both Bombay and further north in the Ganges plain.  Obliques with shorter focal length lenses were requested to document this event.  A striking ISS/CEO image of the southern Himalayas of 11/27 showed noticeable aerosols even in a near vertical view of mountain valleys),Swampland complex, Chad(tropical rainfall generates vast swamps in Chad’s moister south.  A mapping strip for ~1.5 minutes was requested at nadir and just right of track to help map margins of these features which change through time.  Modern margins are of interest to ecologists, ancient margins to geologists since they indicate on-continent sedimentation of vast proportions [handheld imagery is ideal for determining rivers responsible for this sedimentation]),Sahel biomass burning(Dynamic event.  Burning still reported in West Africa.  Pointing left and right), La Paz, Bolivia (pointing right for this city which has the distinction of spanning the greatest altitudinal zone of any in the world–from the flat desert with the airport on the high Andes, down thousands of feet into a major [and warmer] east-facing canyon), Tropical depression, Caribbean(Dynamic event: Out-of-season tropical depression #20 is strengthening in the northern Caribbean, and was predicted to become Tropical Storm Odette when ISS flew over, centered just south of Haiti where storm watches have been posted.  Looking left about 3.5 degrees for a synoptic view of this well organized, unusual storm),Tuamotu-Austral Islands (looking left and right for fringing coral reefs around atolls in this remote island chain), Wichita River(images of the Wichita-Red River confluence were requested to document soil erosion in the area),Spring plankton, New Zealand (Dynamic event: Major bloom underway along most coasts of New Zealand, especially under track on both coasts of North Island), Palmerston Island, Pacific (pointing right of track for this small island.  Detailed image requested for the coral reef atlas.  When images are successfully acquired, sites are removed from the science target list), and Great Salt Lake, Utah (images to show the status of lake levels were requested for an ongoing study).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:00pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 20A.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.3; ppO2 (mmHg) — 158.3; ppCO2 (mmHg) — 3.7.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.3.
  • Node:  Pressure (mmHg) — 752.16; temperature (deg C) — 21.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 754.31; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 754.51; temperature (deg C) — 22.8; shell heater temp (deg C) — 23.3, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 21.3
  • PMA-2:  Shell heater temp (deg C) — 18.5

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in fixed (dual-angle) position (non-suntracking). 
  • SM batteries:  Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting; all other batteries (7) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available:3703 kg (8164 lb) as of 12/4/03 [SM(755) + FGB(2596) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH  XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:35am EST [= epoch]):

  • Mean altitude — 372.4 km
  • Apogee — 376.5 km
  • Perigee — 368.3 km
  • Period — 92.0 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006032
  • Orbits per 24-hr. day — 15.65
  • Solar Beta Angle — -25.8 deg (magnitude increasing)
  • Mean altitude loss in last 24 hours — 150 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28777
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see

SpaceRef staff editor.